OpenStudy (anonymous):

Find the distance between the points given: (2, 2) and (5, 5) A. 3 B. 3 rad 2 C. 7 rad 2

OpenStudy (anonymous):

Lets use the distant between two points formula. \[d = \sqrt{(x2-x1)^{2}+(y2-y1)^{2}}\] now (2,2) and (5,5) x1,y1 x2,y2 so try to substitute those values into the formula and you should get a value and that is your answer.

OpenStudy (anonymous):

attempt to do it yourself and I will tell u if u achieve the right answer.

OpenStudy (anonymous):

i did but well i got \[\sqrt{50}\] nd wen i reduced it i got 7.07... , but none of those are there 0.o

OpenStudy (anonymous):

\[Distance = \sqrt{(y_2 - y_1)^2 + (x_2 - x_1)^2}\]

OpenStudy (anonymous):

\[D = \sqrt{(5 - 2)^2 + (5-2)^2} = \sqrt{9 + 9}\] Can you solve now??

OpenStudy (anonymous):

ohh crapp i was doing da wrong problem over here lol hold up

OpenStudy (anonymous):

Do you know how to solve \[\sqrt{18}\] ??

OpenStudy (anonymous):

\[\sqrt{((5-2)^2+(5-2)^2}\] \[=3\sqrt{2}\]

OpenStudy (anonymous):

@JayDS let felly solve this on her own..

OpenStudy (anonymous):

If you can use the calculator you can just type it in.

OpenStudy (anonymous):

yep I know, I was doing that but you butted in.

OpenStudy (anonymous):

That does not mean that you will solve it on your own...

OpenStudy (anonymous):

how is it \[\sqrt[3]{2}\] tho?? bcus \[\sqrt{18}\] reduces to 4.24....

OpenStudy (anonymous):

Firstly make the factors of 18... Can you do that??

OpenStudy (anonymous):

yeahh

OpenStudy (anonymous):

What are the factors that makes 18???

OpenStudy (anonymous):

1,2,3,6,9,18

OpenStudy (anonymous):

Yes, now from these 18 can be made from which numbers??

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

See 20 has factors: 1, 2 , 4, 5, 10 So, from these numbers I want to form 20.. How can I get that?? 20 = 4*5 Or 20 = 2*2*5 Getting my point?? If yes, then what are the factors that 18 comprises of??

OpenStudy (anonymous):

oh, 1*18, 2*9, 3*6, 6*3, 9*2

OpenStudy (anonymous):

See you said: 18 = 2*9 or 9*2 I say 18 = 3*3*2 Am I right?? You can use the factors twice and thrice also.. Getting?

OpenStudy (anonymous):

oooohh, ii get it now! thank youu !!! :)

OpenStudy (anonymous):

Okay welcome dear...